OpenStudy (anonymous):

Tell whether the system has one solution, infinitely solution, or no solution. 1/3y = 7/3x + 5/3 x-3y = 5

OpenStudy (anonymous):

The first solution is \[\left(\frac{1}{10} \left(21-\sqrt{1141}\right) ,\frac{1}{30} \left(-29-\sqrt{1141}\right )\right) \] The second solution is \[\left(\frac{1}{10} \left(21+\sqrt{1141}\right) ,\frac{1}{30} \left(\sqrt{1141}-29\right) \right) \] Try to find them.